Solve the following Bernoulli’s equation: 𝑑𝑥 /𝑑𝑦 = 𝑥2 + 2𝑥𝑦 + 𝑦2


1
Expert's answer
2021-01-07T14:32:54-0500

"\displaystyle

\frac{\mathrm{d}x}{\mathrm{d}y} = x^2 + 2xy + y^2\\

\frac{\mathrm{d}x}{\mathrm{d}y} = (x + y)^2\\

\frac{\mathrm{d}y}{\mathrm{d}x} = \frac{1}{(x + y)^2}\\

(x + y)^2\frac{\mathrm{d}y}{\mathrm{d}x} = 1\\


\frac{(x + y)^2}{1 + (x + y)^2}\frac{\mathrm{d}y}{\mathrm{d}x} = \frac{1}{1 + (x + y)^2}\\



\frac{1 + (x + y)^2 - 1}{1 + (x + y)^2}\frac{\mathrm{d}y}{\mathrm{d}x} = \frac{1}{1 + (x + y)^2}\\

\frac{\mathrm{d}y}{\mathrm{d}x} - \frac{1}{1 + (x + y)^2}\frac{\mathrm{d}y}{\mathrm{d}x} = \frac{1}{1 + (x + y)^2}\\


\frac{\mathrm{d}y}{\mathrm{d}x} = \frac{1}{1 + (x + y)^2}\left(1 + \frac{\mathrm{d}y}{\mathrm{d}x}\right)\\


\frac{\mathrm{d}y}{\mathrm{d}x} = \frac{1}{1 + (x + y)^2}\cdot\frac{\mathrm{d}}{\mathrm{d}x}\left(x + y\right)\\

\frac{\mathrm{d}y}{\mathrm{d}x} = \frac{\mathrm{d}}{\mathrm{d}x}\left(\arctan(x + y)\right)\\


\int\frac{\mathrm{d}y}{\mathrm{d}x} \cdot \mathrm{d}x= \int \frac{\mathrm{d}}{\mathrm{d}x}\left(\arctan(x + y)\right) \cdot \mathrm{d}x\\


y = \arctan(x + y) + C"